OpenStudy (anonymous):

Solve: |2n - 3| = 5 A. {-1, 4} B. {1, 4} C. {-1, 1} D. {1, -4}

hartnn (hartnn):

Hi :) if there's, |a| = b, we write its as a=b or a = -b so what will you write with |2n-3| =5 ?

OpenStudy (anonymous):

what do you mean ?

hartnn (hartnn):

like for example if we had |9x+7| = 4 we would split it into 2 equations, 9x+7=4 or 9x+7 =-4 how would you split your equation?

OpenStudy (anonymous):

2n - 3 = 5 or -5 when 2n-3 = 5 n=4 when 2n-3 = -5 n = -1

OpenStudy (anonymous):

as |-5|=5 |5|=5

OpenStudy (anonymous):

|2n-3| =5 |2n-3| =-5

OpenStudy (anonymous):

no the |2n-3| must be positive

OpenStudy (anonymous):

oh :3

hartnn (hartnn):

since you have already split the equation, there's no need to use |...| absolute sign 2n-3 =5 or 2n-3 =-5 solve these 2 independently

OpenStudy (anonymous):

-1,4?

hartnn (hartnn):

yes!

OpenStudy (anonymous):

thank you guys so much for your time and help !!

hartnn (hartnn):

you're most welcome ^_^